# Approximation in Hölder spaces

### v2

@@ 2. Bounded support approximation
by $\omega$ being non-decreasing. This concludes the proof of $\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)\subset\overline{C^{\infty}_{c}(\mathbb{R}^{n},Y)}^{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}.$ For the reverse inclusion ``","```$\supset$",` ` ⊃ " , simply notice that by the Mean Value Inequality, $C^{\infty}_{c}(\mathbb{R}^{n},Y)\subset\operatorname{Lip}_{\operatorname{bs}}(\mathbb{R}^{n},Y)$. We remind that this inequality holds for smooth $Y$-valued functions, as a consequence of the Hahn-Banach theorem on $Y.$ Also, by (3) there holds that $\operatorname{Lip}_{\operatorname{bs}}(\mathbb{R}^{n},Y)\subset\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y).$ As $\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)$ is closed with respect to $\dot{C}^{0,\omega}$ limits, we are done. ∎
Proving approximations in a more general normed space $X$ and $Y=\mathbb{R}$ (or $Y=\mathbb{C}$) is a much more delicate task and this is the content of the next Section 3. We finish this section by establishing the connection between pointwise and mean oscillation type conditions.
@@
We first prove the norms equivalent. So fix an arbitrary cube $Q\in\mathcal{Q}$ and estimate
@@
thus we obtain $\|f\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}\gtrsim_{n,C_{\operatorname{db}}}\|f\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}.$ For the other direction, let $x,z\in\mathbb{R}^{n}$ be arbitrary and let $Q_{0}\subset\mathbb{R}^{n}$ be a cube containing both $x$ and $y$ such that $\ell(Q_{0})=|x-y|.$ Consider the dyadic descendants of $Q_{0}$ (achieved by iteratively halving the sides) and for every $k\in\mathbb{N}\cup\{0\}$ let $Q_{k}(x)$ be the descendant of $Q_{0}$ of sidelength $2^{-k}\ell(Q_{0})$ that contains the point $x.$ Similarly, we define $Q_{k}(y)$ for each $k.$ By the continuity of $f:\mathbb{R}^{n}\to Y,$ we write
@@
By the condition $\omega(\infty)=\infty$, we choose $N$ large enough and bound the bracketed term from above by $\leq 2\varepsilon[\omega]_{*}.$ Repeating the same proof with $y$ in place of $x$, we control the right sum on the line (33) and thus obtain $|f(x)-f(y)|\leq 4\varepsilon[\omega]_{*}\omega(|x-y|),$ that is, $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$ Then, we show that $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y)\subset\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$ Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y)$, $\varepsilon>0,$ and $R>0$ be such that $\operatorname{osc}^{\omega}_{\delta}(f)\leq\varepsilon,$ if $\delta\geq R.$ Let $M\gg R$ be a large constant, which we will specify later, and $Q\in\mathcal{Q}$ a cube such that $\ell(Q)\geq M.$ For every $y\in Q,$ denote by $Q_{R}(y)$ the cube centered at $y$ and of $\operatorname{diam}(Q_{R}(y))=R.$ Then, we have
@@ 3.1. Lipschitz and smooth approximations
In a real normed space $X$, we consider real-valued functions ``$f"$ of the class $\dot{\operatorname{VC}}^{0,\omega}(X,\mathbb{R})$, which we abbreviate by $\dot{\operatorname{VC}}^{0,\omega}(X).$ By Theorem 1.4 these functions ``$f"$ can be approximated by functions, say ``$g",$ of class $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ with bounded support, in the $\dot{C}^{0,\omega}(X)$ seminorm. Our goal is to further approximate these functions ``$g"$ by Lipschitz functions (whose regularity will vary depending on the smoothness properties of $X$) with bounded support. In approximating ``$g"$ a crucial step is to reduce the problem to the problem of approximating Lipschitz functions by smooth Lipschitz functions with good control on the Lipschitz constants. The main results of this subsection are Theorem 3.3, valid for arbitrary normed spaces, and Theorem 3.5, valid for separable normed spaces $X$ with a fixed degree of smoothness. The proofs of these two Theorems 3.3 and 3.5 employ several technical lemma, which will be reused in the later Subsections 3.2 and 3.3.
@@
converges uniformly to $f$ in $X$ and each $f_{n}$ is $n$-Lipschitz; for a proof see e.g. [HajJoh14, p. 408]. Now, let us see that $\|f-f_{n}\|_{\dot{C}^{0,\omega}(X)}\to 0,$ as $n\to\infty.$ Indeed, given $\varepsilon>0$ there exists $\delta>0$ so that $\|x-z\|\leq\delta$ implies $|f(x)-f(z)|\leq\varepsilon\omega(\|x-z\|).$ Let us take $N\in\mathbb{N}$ large enough so that $\sup_{X}|f-f_{n}|\leq\varepsilon\omega(\delta)$ for each $n\geq N.$ Using that $\omega$ is non-decreasing, this gives the estimate
@@
The same argument swapping $x$ and $z,$gives $|(f-f_{n})(x)-(f-f_{n})(z)|\leq 2\varepsilon\omega(\|x-z\|),$ for every $n\in\mathbb{N}.$ We conclude that $\|f-f_{n}\|_{\dot{C}^{0,\omega}(X)}\leq 2\varepsilon,$ for $n\geq N.$ For the assertion concerning the boundedness of the supports, we may assume that $f(y)=0,$ whenever $\|y\|\geq R,$ for some $R>0.$ We will localize the infimum defining $f_{n}.$ For each $y\in X,$ it follows from the definition of $f_{n}$, and the fact that $f\in\dot{C}^{0,\omega_{f}}(X),$ that
@@
For the converse implication ``"``"`$\subset$"` ` ⊂ ", it is enough to observe that $\mathrm{Lip}_{\operatorname{bs}}(X)\subset\dot{\operatorname{VC}}^{0,\omega}(X)$ and that $\dot{\operatorname{VC}}^{0,\omega}(X)$ is a closed subspace of $\dot{C}^{0,\omega}(X).$
@@
Let $k\in\mathbb{N}\cup\{\infty\}$ be fixed and let $X$ be a normed space that admits a $C^{k}\cap\operatorname{Lip}$ bump. Let $g\in\operatorname{Lip}(X)$ and $(g_{n})_{n}\subset C^{k}(X)\cap\operatorname{Lip}(X)$ be a sequence of functions converging uniformly to $g$ on $X,$ and with $\limsup_{n}\operatorname{Lip}(g_{n})<\infty.$ Then, if $g$ has bounded support, there exists a sequence $(h_{n})_{n}\subset C^{k}(X)\cap\operatorname{Lip}(X)$ with bounded supports that converge to $g$ uniformly (on $X$) and moreover $\limsup_{n}\operatorname{Lip}(h_{n})<\infty.$
@@
By the properties of $\theta$ and $\psi,$ it is immediate that $\varphi\in C^{k}(X)\cap\operatorname{Lip}(X)$, and that $\varphi$ takes values in $[0,1].$ Also, if $\|x\|\leq 2R,$ one has $\psi(x)\leq 2R$ by (36), and thus $\varphi(x)=\theta(\psi(x))=1.$ Similary, we deduce that $\varphi(x)=\theta(\psi(x))=0$ for those $x\in X$ such that $\|x\|\geq 4R/a=:\lambda.$ Now, by assumption we can find a sequence $(g_{n})_{n}\subset C^{k}(X)\cap\operatorname{Lip}(X)$ converging uniformly to $g$ on $X,$ and, after passing to a subsequence, with the property $\sup_{n}\operatorname{Lip}(g_{n})<\infty.$ Define $h_{n}:=\varphi g_{n}$ for each $n\in\mathbb{N},$ where $\varphi$ is as in (37). Since $\varphi$ vanishes outside the ball $B(0,4R/a),$ the function $g_{n}$ has bounded support. Also,
@@
Thus we have shown the inclusion ``"."```$\supset$".` ` ⊃ " . For the converse inclusion ``"``"`$\subset$"` ` ⊂ ", it is enough to observe again (as we did at the end of the proof of Theorem 3.3) that $\mathrm{Lip}_{\operatorname{bs}}(X)\subset\dot{\operatorname{VC}}^{0,\omega}(X)$ and that $\dot{\operatorname{VC}}^{0,\omega}(X)$ is a closed subspace of ${\dot{C}^{0,\omega}(X)}.$
@@
Both statements (i) and (ii) follow immediately from Theorem 3.5, since a Hilbert space always admits a $C^{\infty}\cap\operatorname{Lip}$ bump. For the sake of completeness, we exhibit an elementary construction of such a bump. The function $\psi:X\to\mathbb{R}$ given by $x\mapsto\psi(x)=\|x\|^{2}$ is of class $C^{\infty}(X)$ with the Fréchet derivative $D\psi(x):X\to\mathbb{R},$ at every $x\in X$, given by $D\psi(x)(v)=\langle 2x,v\rangle$ for all $v\in X.$ Take a function $\theta:\mathbb{R}\to[0,1]$ of class $C^{\infty}(\mathbb{R})\cap\operatorname{Lip}(\mathbb{R})$ with $\theta(t)=1$ for all $t\leq 1/2$ and $\theta(t)=0$ for all $t\geq 1.$ Then, the map $\varphi=\theta\circ\psi:X\to[0,1]$ is of class $C^{\infty}(X)$ with $\varphi(x)=1$ for all $\|x\|\leq 1/2$ and $\varphi(x)=0$ for all $\|x\|\geq 1.$ Also, the Fréchet derivative of $\varphi$ satisfies
@@ 3.2. Approximation in super-reflexive spaces
Let $g:X\to\mathbb{R}$ be $L$-Lipschitz, and $\varepsilon>0.$ We will now apply part (1) for a suitable sequence of sets to obtain the desired approximation. In order to do so, we use the same construction as in [HajJoh10, Proposition 1], replacing the $C^{1}$-separating functions from there with those we just obtained in part (1). Define $\tilde{g}(x)=\varepsilon^{-1}g(\varepsilon x/L),$ for all $x\in X,$ and note that $\tilde{g}$ is $1$-Lipschitz. Now, we define the set $S_{n}=\{x\in X\,:\,\tilde{g}(x)\geq n\},$ for each integer $n\in\mathbb{Z}.$ Note that $S_{n+1}\subset S_{n}$, and also, because $\tilde{g}$ is $1$-Lipschitz, one has that $d(X\setminus S_{n},S_{n+1})\geq 1$ for every $n\in\mathbb{Z}.$ Now, for each subset $S_{n},$ let $h_{S_{n}}:X\to[0,1]$ be the function from (1). Define now $h_{n}=1-h_{S_{n+1}}$ for every $n$, and also
@@
The relation (41) will allow us to differentiate the sums defining $h$ term by term locally around every $x\in X.$ Let us now prove that $h\in C^{1,\alpha}(X)$ with $\|Dh\|_{\dot{C}^{0,\alpha}(X,X^{*})}\leq 2\kappa.$ Recalling that by definition $Dh_{S_{n+1}}(x)=0$ whenever $x\in S_{n+1}\cup\{x\in X\,:\,d(x,S_{n+1})\geq 1\}$, it is immediate from the definition $h_{n}:=1-h_{S_{n+1}}$ that
@@
This proves the claim (43). Now given $x,y\in X,$ let $m\in\mathbb{Z}$ and $l\in\mathbb{N}\cup\{0\}$ be so that $x\in S_{m}\setminus S_{m+1}$ and $y\in S_{m+l}\setminus S_{m+l+1},$ and we show the $\alpha$-Hölder estimate for $\|Dh(x)-Dh(y)\|_{*}.$ In the case $l=0,$ we have $x,y\in S_{m}\setminus S_{m+1},$ and so it suffices to apply (43) and the fact that $\|Dh_{m}\|_{\dot{C}^{0,\alpha}(X,X^{*})}\leq\kappa.$ Assume now $l\geq 1$. By the connectedness of $[x,y]$ and $x\not\in S_{m+l}$ and $y\in S_{m+l}$, the segment $[x,y]$ must intersect the boundary of $S_{m+l}.$ Let $z\in[x,y]\cap S_{m+l}\cap\overline{X\setminus S_{m+l}}$. In particular, $z\in\overline{X\setminus S_{m+l}},$ and by (42), this implies $Dh_{m+l}(z)=0.$ Also, observe that $y\in S_{m+l}\subset S_{m+1},$ and so $Dh_{m}(y)=0,$ again by (42). Using first (43), then $Dh_{m}(y)=Dh_{m+l}(z)=0$, and finally that $\|Dh_{n}\|_{\dot{C}^{0,\alpha}(X,X^{*})}\leq\kappa$ for all $n\in\mathbb{Z},$ we conclude that
@@
and let us prove that both terms define an $\alpha$-Hölder function. For the first function $x\mapsto g_{n}(x)D\varphi(x)$, let us observe that, for $n$ large enough, one has $\sup_{X}|g_{n}|\leq 1+\sup_{X}|g|,$ as $g_{n}$ converges uniformly on $X$ to $g.$ Hence, by Remark 3.9, $g_{n}$ is $\alpha$-Hölder continuous on $X.$ Then, we write
@@
All the factors that multiply the term $\|x-z\|^{\alpha}$ are finite by the properties of $\varphi$ and $g_{n}.$ As concerns the second function $x\mapsto\varphi(x)Dg_{n}(x)$, its $\alpha$-Hölder continuity is verified in a very similar way, this time using Remark 3.9 for $\varphi$ and that $Dg_{n}$ is bounded (as $g_{n}$ is Lipschitz) and $\alpha$-Hölder continuous. We conclude that $(h_{n})_{n}\to g$ uniformly, with $(h_{n})_{n}\in C^{1,\alpha}_{\operatorname{bs}}(X)$ and $\limsup_{n}\operatorname{Lip}(h_{n})<\infty.$ By Lemma 3.2, we have $\|h_{n}-g\|_{\dot{C}^{0,\omega}(X)}\to 0$ as well. Therefore, we can find $h\in C^{1,\alpha}_{\operatorname{bs}}(X)$ with $\|h-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon,$ and thus $\|h-f\|_{\dot{C}^{0,\omega}(X)}\leq 2\varepsilon.$ We have shown the inclusion $\dot{\operatorname{VC}}^{0,\omega}(X)\subset\overline{C^{1,\alpha}_{\operatorname{bs}}(X)}^{\dot{C}^{0,\omega}(X)}.$ For the reverse inclusion, observe that if $f\in C^{1,\alpha}_{\operatorname{bs}}(X),$ then $Df:X\to X^{*}$ is $\alpha$-Hölder and with bounded support, and so $Df$ is bounded in $X.$ Because Lipschitz functions with bounded support are contained in $\dot{\operatorname{VC}}^{0,\omega}(X),$ which is closed under limits with respect to the $\dot{C}^{0,\omega}(X)$ seminorm, the reverse inclusion holds. ∎

### v3

@@ 2. Bounded support approximation
by $\omega$ being non-decreasing. This concludes the proof of $\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)\subset\overline{C^{\infty}_{c}(\mathbb{R}^{n},Y)}^{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}.$
For the reverse inclusion ``","```$\supset$",` ` ⊃ " , simply notice that by the Mean Value Inequality, $C^{\infty}_{c}(\mathbb{R}^{n},Y)\subset\operatorname{Lip}_{\operatorname{bs}}(\mathbb{R}^{n},Y)$. We remind that this inequality holds for smooth $Y$-valued functions, as a consequence of the Hahn-Banach theorem on $Y.$ Also, by (3) there holds that $\operatorname{Lip}_{\operatorname{bs}}(\mathbb{R}^{n},Y)\subset\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y).$ As $\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)$ is closed with respect to $\dot{C}^{0,\omega}$ limits, by Remark 2.1, we are done. ∎
Next, we record a quantitative version of Theorem 1.6 to be used in the proof of Theorem 1.11 in [MarOik24]. While Corollary 2.5 below is one the key elements in applications to the bi-commutator in [MarOik24], it plays no further role in the present article.
Definition 2.4.
For a subset $G$ of $\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y),$ we write $G\subset_{u}\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)$ (with ``$u"$ for uniformly) provided that $G$ is equibounded in the $\dot{C}^{0,\omega}$ seminorm, and that for all $\varepsilon>0,$ there exists $t>0$ such that if $|x-y|<t$ or $|x|>t^{-1}$ or $|y|>t^{-1},$ then
$\sup_{g\in G}\frac{|g(x)-g(y)|}{\omega(|x-y|)}<\varepsilon.$
By an inspection of the above proofs, the reader convinces themselves that the following is true.
Corollary 2.5.
Let $Y$ be a Banach space and the modulus $\omega$ satisfy (2) and (4). Let $G\subset_{u}\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y).$ Then,
  • for all $L>0$ we have $G\circ\tau_{L}\subset_{u}\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y),$
  • for all $L,L^{\prime}>0$ we have $(G\circ\tau_{L})*\eta_{L^{\prime}}\subset_{u}\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y),$
  • for all $M=M(\varepsilon)>0$ sufficiently large
    $\sup_{f\in G}\|f-f\circ\tau_{M}\|_{\dot{\operatorname{C}}^{0,\omega}(\mathbb{R}^{n},Y)}\leq\varepsilon,$ (26)
  • for all $K=K(M,\varepsilon)>0$ sufficiently large
    $\sup_{f\in G}\|f\circ\tau_{M}-(f\circ\tau_{M})*\eta_{1/K}\|_{\dot{\operatorname{C}}^{0,\omega}(\mathbb{R}^{n},Y)}\leq\varepsilon.$ (27)
Proving approximations with a more general normed space $X$ and $Y=\mathbb{R}$ (or $Y=\mathbb{C}$) is a delicate task and this is the content of the next Section 3. But before that, we finish this section by proving the connection between pointwise and mean oscillation type conditions.
@@
We first prove that the norms are equivalent. So fix an arbitrary cube $Q\in\mathcal{Q}$ and estimate
@@
thus we obtain $\|f\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}\gtrsim_{n,C_{\operatorname{db}}}\|f\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}.$
For the other direction, let $x,z\in\mathbb{R}^{n}$ be arbitrary and let $Q_{0}\subset\mathbb{R}^{n}$ be a cube containing both $x$ and $y$ such that $\ell(Q_{0})=|x-y|.$ Consider the dyadic descendants of $Q_{0}$ (achieved by iteratively halving the sides) and for every $k\in\mathbb{N}\cup\{0\}$ let $Q_{k}(x)$ be the descendant of $Q_{0}$ of sidelength $2^{-k}\ell(Q_{0})$ that contains the point $x.$ Similarly, we define $Q_{k}(y)$ for each $k.$ By the continuity of $f:\mathbb{R}^{n}\to Y,$ we write
@@
By the condition $\omega(\infty)=\infty$, we choose $N$ large enough and bound the bracketed term from above by $\leq 2\varepsilon[\omega]_{*}.$ Repeating the same proof with $y$ in place of $x$, we control the right sum on the line (29) and thus obtain $|f(x)-f(y)|\leq 4\varepsilon[\omega]_{*}\omega(|x-y|),$ that is, $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$
Then, we show that $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y)\subset\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$ Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y)$, $\varepsilon>0,$ and $R>0$ be such that $\operatorname{osc}^{\omega}_{\delta}(f)\leq\varepsilon,$ if $\delta\geq R.$ Let $M\gg R$ be a large constant, which we will specify later, and $Q\in\mathcal{Q}$ a cube such that $\ell(Q)\geq M.$ For every $y\in Q,$ denote by $Q_{R}(y)$ the cube centered at $y$ and of $\operatorname{diam}(Q_{R}(y))=R.$ Then, we have
@@ 3.1. Lipschitz and smooth approximations
In a real normed space $X$, we consider real-valued functions ``$f"$ of the class $\dot{\operatorname{VC}}^{0,\omega}(X,\mathbb{R})$, the last abbreviated by $\dot{\operatorname{VC}}^{0,\omega}(X).$ By Theorem 1.2 these functions ``$f"$ can be approximated by functions, say ``$g",$ of class $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ with bounded support, in the $\dot{C}^{0,\omega}(X)$ seminorm. Our goal is to further approximate these functions ``$g"$ by Lipschitz functions (whose regularity will vary depending on the smoothness properties of $X$) with bounded support. In approximating ``$g"$ a crucial step is to reduce the problem to the problem of approximating Lipschitz functions by smooth Lipschitz functions with good control on the Lipschitz constants. The main results of this subsection are Theorem 3.3, valid for arbitrary normed spaces, and Theorem 3.5, valid for separable normed spaces $X$ with a fixed degree of smoothness. The proofs of these two Theorems 3.3 and 3.5 employ several technical lemma, which will be reused in the later Subsections 3.2 and 3.3.
@@
converges uniformly to $f$ in $X$ and each $f_{n}$ is $n$-Lipschitz; for a proof see e.g. [HajJoh14, p. 408].
Now, let us see that $\|f-f_{n}\|_{\dot{C}^{0,\omega}(X)}\to 0,$ as $n\to\infty.$ Indeed, given $\varepsilon>0$ there exists $\delta>0$ so that $\|x-z\|\leq\delta$ implies $|f(x)-f(z)|\leq\varepsilon\omega(\|x-z\|).$ Let us take $N\in\mathbb{N}$ large enough so that $\sup_{X}|f-f_{n}|\leq\varepsilon\omega(\delta)$ for each $n\geq N.$ Using that $\omega$ is non-decreasing, this gives the estimate
@@
The same argument swapping $x$ and $z,$gives $|(f-f_{n})(x)-(f-f_{n})(z)|\leq 2\varepsilon\omega(\|x-z\|),$ for every $n\in\mathbb{N}.$ We conclude that $\|f-f_{n}\|_{\dot{C}^{0,\omega}(X)}\leq 2\varepsilon,$ for $n\geq N.$
For the assertion concerning the boundedness of the supports, we may assume that $f(y)=0,$ whenever $\|y\|\geq R,$ for some $R>0.$ We will localize the infimum defining $f_{n}.$ For each $y\in X,$ it follows from the definition of $f_{n}$, and the fact that $f\in\dot{C}^{0,\omega_{f}}(X),$ that
@@
For the converse implication ``"``"`$\subset$"` ` ⊂ ", it is enough to observe that $\mathrm{Lip}_{\operatorname{bs}}(X)\subset\dot{\operatorname{VC}}^{0,\omega}(X)$ and that $\dot{\operatorname{VC}}^{0,\omega}(X)$ is a closed subspace of $\dot{C}^{0,\omega}(X);$ see Remark 2.1. ∎
@@
Let $k\in\mathbb{N}\cup\{\infty\}$ be fixed and let $X$ be a normed space that admits a $C^{k}\cap\operatorname{Lip}$ bump. Let $g\in\operatorname{Lip}(X)$ and $(g_{n})_{n}\subset C^{k}(X)\cap\operatorname{Lip}(X)$ be a sequence of functions converging uniformly to $g$ on $X,$ and with $\limsup_{n}\operatorname{Lip}(g_{n})<\infty.$
Then, if $g$ has bounded support, there exists a sequence $(h_{n})_{n}\subset C^{k}(X)\cap\operatorname{Lip}(X)$ with bounded supports that converge to $g$ uniformly (on $X$) and moreover $\limsup_{n}\operatorname{Lip}(h_{n})<\infty.$
@@
By the properties of $\theta$ and $\psi,$ it is immediate that $\varphi\in C^{k}(X)\cap\operatorname{Lip}(X)$, and that $\varphi$ takes values in $[0,1].$ Also, if $\|x\|\leq 2R,$ one has $\psi(x)\leq 2R$ by (32), and thus $\varphi(x)=\theta(\psi(x))=1.$ Similary, we deduce that $\varphi(x)=\theta(\psi(x))=0$ for those $x\in X$ such that $\|x\|\geq 4R/a=:\lambda.$
Now, by assumption we can find a sequence $(g_{n})_{n}\subset C^{k}(X)\cap\operatorname{Lip}(X)$ converging uniformly to $g$ on $X,$ and, after passing to a subsequence, with the property $\sup_{n}\operatorname{Lip}(g_{n})<\infty.$ Define $h_{n}:=\varphi g_{n}$ for each $n\in\mathbb{N},$ where $\varphi$ is as in (33). Since $\varphi$ vanishes outside the ball $B(0,4R/a),$ the function $g_{n}$ has bounded support. Also,
@@
Thus we have shown the inclusion ``"."```$\supset$".` ` ⊃ " . For the converse inclusion ``"``"`$\subset$"` ` ⊂ ", it is enough to observe again (as we did at the end of the proof of Theorem 3.3) that $\mathrm{Lip}_{\operatorname{bs}}(X)\subset\dot{\operatorname{VC}}^{0,\omega}(X)$ and that $\dot{\operatorname{VC}}^{0,\omega}(X)$ is a closed subspace of ${\dot{C}^{0,\omega}(X)},$ again, by Remark 2.1.
@@
Both statements (i) and (ii) follow immediately from Theorem 3.5, since a Hilbert space always admits a $C^{\infty}\cap\operatorname{Lip}$ bump. For the sake of completeness, we exhibit an elementary construction of such a bump.
The function $\psi:X\to\mathbb{R}$ given by $x\mapsto\psi(x)=\|x\|^{2}$ is of class $C^{\infty}(X)$ with the Fréchet derivative $D\psi(x):X\to\mathbb{R},$ at every $x\in X$, given by $D\psi(x)(v)=\langle 2x,v\rangle$ for all $v\in X.$ Take a function $\theta:\mathbb{R}\to[0,1]$ of class $C^{\infty}(\mathbb{R})\cap\operatorname{Lip}(\mathbb{R})$ with $\theta(t)=1$ for all $t\leq 1/2$ and $\theta(t)=0$ for all $t\geq 1.$ Then, the map $\varphi=\theta\circ\psi:X\to[0,1]$ is of class $C^{\infty}(X)$ with $\varphi(x)=1$ for all $\|x\|\leq 1/2$ and $\varphi(x)=0$ for all $\|x\|\geq 1.$ Also, the Fréchet derivative of $\varphi$ satisfies
@@ 3.2. Approximation in super-reflexive spaces
Let $g:X\to\mathbb{R}$ be $L$-Lipschitz, and $\varepsilon>0.$ We will now apply part (1) for a suitable sequence of sets to obtain the desired approximation. In order to do so, we use the same construction as in [HajJoh10, Proposition 1], replacing the $C^{1}$-separating functions from there with those we just obtained in part (1).
Define $\tilde{g}(x)=\varepsilon^{-1}g(\varepsilon x/L),$ for all $x\in X,$ and note that $\tilde{g}$ is $1$-Lipschitz. Now, we define the set $S_{n}=\{x\in X\,:\,\tilde{g}(x)\geq n\},$ for each integer $n\in\mathbb{Z}.$ Note that $S_{n+1}\subset S_{n}$, and also, because $\tilde{g}$ is $1$-Lipschitz, one has that $d(X\setminus S_{n},S_{n+1})\geq 1$ for every $n\in\mathbb{Z}.$ Now, for each subset $S_{n},$ let $h_{S_{n}}:X\to[0,1]$ be the function from (1). Define now $h_{n}=1-h_{S_{n+1}}$ for every $n$, and also
@@
The relation (37) will allow us to differentiate the sums defining $h$ term by term locally around every $x\in X.$
Let us now prove that $h\in C^{1,\alpha}(X)$ with $\|Dh\|_{\dot{C}^{0,\alpha}(X,X^{*})}\leq 2\kappa.$ Recalling that by definition $Dh_{S_{n+1}}(x)=0$ whenever $x\in S_{n+1}\cup\{x\in X\,:\,d(x,S_{n+1})\geq 1\}$, it is immediate from the definition $h_{n}:=1-h_{S_{n+1}}$ that
@@
This proves the claim (39).
Now given $x,y\in X,$ let $m\in\mathbb{Z}$ and $l\in\mathbb{N}\cup\{0\}$ be so that $x\in S_{m}\setminus S_{m+1}$ and $y\in S_{m+l}\setminus S_{m+l+1},$ and we show the $\alpha$-Hölder estimate for $\|Dh(x)-Dh(y)\|_{*}.$
In the case $l=0,$ we have $x,y\in S_{m}\setminus S_{m+1},$ and so it suffices to apply (39) and the fact that $\|Dh_{m}\|_{\dot{C}^{0,\alpha}(X,X^{*})}\leq\kappa.$
Assume now $l\geq 1$. By the connectedness of $[x,y]$ and $x\not\in S_{m+l}$ and $y\in S_{m+l}$, the segment $[x,y]$ must intersect the boundary of $S_{m+l}.$ Let $z\in[x,y]\cap S_{m+l}\cap\overline{X\setminus S_{m+l}}$. In particular, $z\in\overline{X\setminus S_{m+l}},$ and by (38), this implies $Dh_{m+l}(z)=0.$ Also, observe that $y\in S_{m+l}\subset S_{m+1},$ and so $Dh_{m}(y)=0,$ again by (38). Using first (39), then $Dh_{m}(y)=Dh_{m+l}(z)=0$, and finally that $\|Dh_{n}\|_{\dot{C}^{0,\alpha}(X,X^{*})}\leq\kappa$ for all $n\in\mathbb{Z},$ we conclude that
@@
and let us prove that both terms define an $\alpha$-Hölder function.
For the first function $x\mapsto g_{n}(x)D\varphi(x)$, let us observe that, for $n$ large enough, one has $\sup_{X}|g_{n}|\leq 1+\sup_{X}|g|,$ as $g_{n}$ converges uniformly on $X$ to $g.$ Hence, by Remark 3.9, $g_{n}$ is $\alpha$-Hölder continuous on $X.$ Then, we write
@@
All the factors that multiply the term $\|x-z\|^{\alpha}$ are finite by the properties of $\varphi$ and $g_{n}.$
As concerns the second function $x\mapsto\varphi(x)Dg_{n}(x)$, its $\alpha$-Hölder continuity is verified in a very similar way, this time using Remark 3.9 for $\varphi$ and that $Dg_{n}$ is bounded (as $g_{n}$ is Lipschitz) and $\alpha$-Hölder continuous.
We conclude that $(h_{n})_{n}\to g$ uniformly, with $(h_{n})_{n}\in C^{1,\alpha}_{\operatorname{bs}}(X)$ and $\limsup_{n}\operatorname{Lip}(h_{n})<\infty.$ By Lemma 3.2, we have $\|h_{n}-g\|_{\dot{C}^{0,\omega}(X)}\to 0$ as well. Therefore, we can find $h\in C^{1,\alpha}_{\operatorname{bs}}(X)$ with $\|h-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon,$ and thus $\|h-f\|_{\dot{C}^{0,\omega}(X)}\leq 2\varepsilon.$
We have shown the inclusion $\dot{\operatorname{VC}}^{0,\omega}(X)\subset\overline{C^{1,\alpha}_{\operatorname{bs}}(X)}^{\dot{C}^{0,\omega}(X)}.$ For the reverse inclusion, observe that if $f\in C^{1,\alpha}_{\operatorname{bs}}(X),$ then $Df:X\to X^{*}$ is $\alpha$-Hölder and with bounded support, and so $Df$ is bounded in $X.$ Because Lipschitz functions with bounded support are contained in $\dot{\operatorname{VC}}^{0,\omega}(X),$ which is closed under limits with respect to the $\dot{C}^{0,\omega}(X)$ seminorm, the reverse inclusion holds. ∎